# Taming Lookup Tables for Efficient Image Retouching

### v1

@@ 3.2.2 Split fully connected layer
where $S$ is the size of LUT, $V$ denotes the possible values in a dimension, and $N$ is the output dimension. When $C>4$, $V=64$ and $N=20$, the LUT is generally beyond 1GB. To avoid substantial memory consumption, we design the split fully connected (SFC) layer. First, we split the input tensor into $K$ groups, and each group contains $L$ values from channel features. Then, we apply a vanilla FC layer to map each 2D feature into a weight of length $N$. These predicted weights are added at the end for weighting the basis LUTs to build a 3D weighted LUT. When we set $L=2$ and $K=C/2$, this manipulation reduces the memory from $(V)^{C}\times N$ to $(C/2)\times(V)^{2}\times N$. Subsequently, when $C=10$, $V=64$ and $N=20$, the LUT size is equal to 400KB, which is orders smaller than the vanilla FC layer. We further compare the performance in Table 7 and observe a minimal performance drop when substituting the FC with an SFC layer.
@@ 3.4 Speeding up inference
Our main goal is to predict a set of weights for weighting the basis LUTs to produce the final 3D LUT. The CNN weight predictor aims to understand the global context, such as brightness, color, and tones of the image, to output content-dependent weights. Therefore, it only needs to work on the downsampled input image to largely lessen the computational cost [8]. Given an input image of any resolution, previous works simply used bilinear interpolation to downsample it to $256\times 256$ for high efficiency. Nevertheless, as the input size increases, the network’s computational workload, or table lookup operation, also escalates. Hence, a reduction in the image resolution can directly speed up inference and provide cost savings. Subsequently, we compare the performance of ours and existing methods. All methods are trained under the $256\times 256$ resolution and tested for various downsampled scales. The results of CLUT are shown in Fig. 3. We observe that at high resolutions, different models consistently performed well. However, as the resolution decreases, the models using a large RF (viz. 3D LUT and CLUT) suffer a sharp performance drop, while ICELUT (with a $1\times 1$ RF) maintains a high level of performance.
@@
Figure 5: Histogram of images at different resolutions.
The reasons for these results can be attributed to two main factors. First, both 3DLUT and CLUT are composed of multiple $3\times 3$ CNN layers, thus inheriting the inductive bias inherent in the large kernels. This results in a significant performance drop when the scales of downsampled images are inconsistent with the training images. Second, all kernels in our method are $1\times 1$, so the interdependence among spatial pixels is decoupled. This enables our method to achieve robust performance at different scales. While it is apparent that downsized images may lose information, when visualizing histograms of different image scales in Fig. 5, we observe that low-resolution images still retain the fundamental color information of the original high-resolution images. This explains that the network can still capture vital image attributes, including brightness, color, and tonal characteristics, and make correct predictions for LUT weights.
4 Experiments
4.1 Datasets
We evaluate the proposed ICELUT using two public datasets: MIT-Adobe FiveK [9] and PPR10K [23]. The MIT-Adobe FiveK dataset is a well-known photo retouching dataset comprising 5,000 RAW images. We follow the common practice established in recent works [8, 6], by selecting the version retouched by expert C as the ground truth. We divide this dataset into 4,500 image pairs for training and 500 for testing. To expedite the training process, we downsized the images to 480p resolution, where the shorter side is resized to 480 pixels. The PPR10K dataset contains an extensive collection of 11,161 high-quality RAW portrait photos. In our experiments, we use all three retouched versions as the ground truth in three separate experiments. Adhering to the official dataset split as in [23], we divide the dataset into 8,875 pairs for training and 2,286 for testing. For the sake of efficiency and due to limited disk space, we perform these experiments using the 360p version of the dataset. Note that the performance data in [23] is not fair in the discussion of runtime and FLOPs since the 3D LUT trained in [23] used a much larger backbone, ResNet18[2] (about 11M). To this point, we have retrained 3D LUT using the original tiny backbone in [8].
@@ 4.2 Implementation details
4.3 Evaluation metrics
We employ three metrics, including peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) [25], and $\Delta E$ to evaluate different methods. $\Delta E$ is a color difference metric defined in the CIELAB color space [24].
@@ 4.4 Quantitative results
4.5 Qualitative results
Despite the very mild PSNR drop compared to other methods, ICELUT achieves uncompromising perceptual quality. Thanks to the pointwise mapping, the image retouched by ICELUT does not exhibit color banding, such as in the cloud in the second row of Fig. 6. Furthermore, the incorporation of RGB channels helps align the images more consistently with their targets, such as the tone in the first row and the exposure in the third row of Fig. 6.
@@ 4.6 Real-time performance comparisons
Table 8: Comparison between different backbones with vanilla FC or split FC.
Backbone Dimension PSNR (w/ FC) PSNR (w/ split FC)
32-64-128-256-512-256-128-64-32 25.33 25.31
16-32-64-128-256-128-64-32-16 25.29 25.14
Table 9: Comparison of one-branch (8-bit input) and two-branch (two 4-bit inputs) networks.
Method PSNR SSIM Channel LUT Size (MB)
Single (8 bits) 25.35 0.919 1,342
Parallel (4 bits) 25.27 0.918 0.20
Table 10: Effect of the number $N$ of basis LUTs. Size denotes the storage of Weight LUTs.
$N$ 1 5 10 15 20 25
PSNR (dB) 23.27 25.16 25.27 25.29 25.30 25.28
Size (KB) 40 200 400 600 800 1,000
@@ 4.7.1 Split FC layer
As described earlier, a split FC layer is employed to save memory while achieving global information fusion effectively. Here, we give the ablation study of this layer. We set the FC output dimension to 20, i.e., 20 weights for weighting 20 basis LUTs. First, we quantize the output from FP32 to INT8 for saving the LUT memory. Note that the quantization error has little impact on the performance. To quantitatively demonstrate the impact on the overall performance and determine the most suitable settings, we first evaluate with the channel numbers $C=\{6,12,18\}$ and split group lengths $L=\{2,3\}$ (for the reason that $L>3$ already leads to an undesirable LUT size). We remark $C=K\times L$, where $K$ is the SFC number. The results are in Table 7. It can be seen that an increase in $K$ leads to improved model performance. To strike a balance between memory size and performance, we choose $C=18$ for subsequent experiments. Compared to a vanilla FC layer, which gathers full channel features for predictions, it is evident that substituting the SFC layer with an FC layer yields only marginal improvements at the expense of a much higher storage cost. Furthermore, we compare our method with the model using pooled features for weight prediction in the last row of Table 7. Obviously, without the global fusion, such as the FC or SFC layers, the features extracted from the pointwise convolution network are insufficient to achieve high performance. On the other hand, the interactions between channels have largely been captured by pointwise convolution (whose wide channel dimension during training will not burden inference), diminishing the significance of FC in establishing global channel connections. It means we could use a broader convolution network as our backbone with the sack of no inference burden. Hence, a weakened version, split FC, is already capable of transforming image information into basis coefficients with a powerful backbone, as shown in Table 8 whose middle layer has the largest dimension.
4.7.2 High-low bit separation
To efficiently construct the Channel LUT for storing the output of each input pixel, we employ two parallel branches to process the MSBs and LSBs independently, resulting in significant LUT memory savings compared to the whole INT8 input. Again, we set the output channel number $C=10$. As depicted in Table 9, while there is a very slight drop in performance, the parallel design substantially reduces memory consumption by a factor of 1500×1500$\times$1500 × vs a single-branch architecture.
4.7.3 Number of basis LUTs
To investigate the effect of the number of basis LUTs $N$, we set $C=10,K=5,L=2$ and $N=\{1,3,5,10,15,20\}$, and compute the sizes of LUTs converted from the SFC layer in Table 10. Notably, when $N$ is small, the expressiveness of color transforms is restricted by the limited representation of basis LUTs. When $N$ is large enough, the performance improvement is only minor.
@@ 4.7.4 Quantization of Weight LUTs for downscaling storage
5 Limitation
Our method generates unsatisfying results on images with large high-contrast smoothing regions, as shown in the Fig. 7. Because average pooling compresses the spatial size into $1\times 1$, the large smooth regions will suffer from global biased tone or brightness from other regions due to the average operation. Nonetheless, such cases occupy only 1.5% of the whole dataset, and are still visually plausible as shown in Fig. 7.
Figure 7: Failure case.
@@ 6 Conclusion
This work has proposed the first-of-its-kind purely LUT-based image enhancer, ICELUT, for extremely low-cost and high-speed image retouching. We reveal that input RGB channels are vital for performance and employ fully pointwise convolution kernels that favor subsequent LUT conversion after training. A novel split fully connected layer is devised to effectively suppress the LUT size without compromising performance. Throughout the development of ICELUT, we discover a small receptive field, in its extreme case of $1\times 1$, can instill robustness to low-resolution inputs for both training and inference. Inspired by this insight, an unprecedented $32\times 32$ downsampled image is leveraged in ICELUT for an extreme speed with minimal performance drop. With the concerted efforts of these novel designs, our purely LUT-based scheme achieves a remarkable 0.4ms (7ms) on GPU (CPU) with near-state-of-the-art performance, and reduces the power consumption to a negligible level compared to all other CNN solutions.

### v2

@@ 3.2.2 Split fully connected layer
where $S$ is the size of LUT, $V$ denotes the possible values in a dimension, and $N$ is the output dimension. When $C>4$, $V=64$ and $N=20$, the LUT is generally beyond 1GB. To avoid substantial memory consumption, we design the split fully connected (SFC) layer. First, we split the input tensor into $K$ groups, and each group contains $L$ values from channel features. Then, we apply a vanilla FC layer to map each 2D feature into a weight of length $N$. These predicted weights are added at the end for weighting the basis LUTs to build a 3D weighted LUT. When we set $L=2$ and $K=C/2$, this manipulation reduces the memory from $(V)^{C}\times N$ to $(C/2)\times(V)^{2}\times N$. Subsequently, when $C=10$, $V=64$ and $N=20$, the LUT size is equal to 400KB, which is orders smaller than the vanilla FC layer.
@@ 3.4 Speeding up inference
Our main goal is to predict a set of weights for weighting the basis LUTs to produce the final 3D LUT. The CNN weight predictor aims to understand the global context, such as brightness, color, and tones of the image, to output content-dependent weights. Therefore, it only needs to work on the downsampled input image to largely lessen the computational cost [23]. Given an input image of any resolution, previous works simply used bilinear interpolation to downsample it to $256\times 256$ for high efficiency. Nevertheless, as the input size increases, the network’s computational workload, or table lookup operation, also escalates. Hence, a reduction in the image resolution can directly speed up inference and provide cost savings. Subsequently, we compare the performance of ours and existing methods. All methods are trained under the $256\times 256$ resolution and tested for various downsampled scales. The results of CLUT are shown in Fig. 3.
@@
We observe that at high resolutions, different models consistently performed well. However, as the resolution decreases, the models using a large RF (viz. 3D LUT and CLUT) suffer a sharp performance drop, while ICELUT (with a $1\times 1$ RF) maintains a high level of performance.
4 Experiments
4.1 Datasets
We evaluate the proposed ICELUT using two public datasets: MIT-Adobe FiveK [3] and PPR10K [13]. The MIT-Adobe FiveK dataset is a well-known photo retouching dataset comprising 5,000 RAW images. We follow the common practice established in recent works [23, 24], by selecting the version retouched by expert C as the ground truth. We divide this dataset into 4,500 image pairs for training and 500 for testing. To expedite the training process, we downsized the images to 480p resolution, where the shorter side is resized to 480 pixels. The PPR10K dataset contains an extensive collection of 11,161 high-quality RAW portrait photos. In our experiments, we use all three retouched versions as the ground truth in three separate experiments. Adhering to the official dataset split as in [13], we divide the dataset into 8,875 pairs for training and 2,286 for testing. For the sake of efficiency and due to limited disk space, we perform these experiments using the 360p version of the dataset. Note that the performance data in [13] are not fair in the discussion of runtime and FLOPs since the 3D LUT trained in [13] used a much larger backbone, ResNet18[7] (about 11M). To this point, we have retrained 3D LUT using the original tiny backbone in [23].
@@ 4.5.1 Split FC layer
As described earlier, a split FC layer is employed to save memory while achieving global information fusion effectively. Here, we give the ablation study of this layer. We set the FC output dimension to 20, i.e., 20 weights for weighting 20 basis LUTs. First, we quantize the output from FP32 to INT8 for saving the LUT memory. Note that the quantization error has little impact on the performance. To quantitatively demonstrate the impact on the overall performance and determine the most suitable settings, we first evaluate with the channel numbers $C=\{6,12,18\}$ and split group lengths $L=\{2,3\}$ (for the reason that $L>3$ already leads to an undesirable LUT size). We remark $C=K\times L$, where $K$ is the SFC number. The results are in Table 7. It can be seen that an increase in $K$ leads to improved model performance. To strike a balance between memory size and performance, we choose $C=18$ for subsequent experiments. Compared to a vanilla FC layer, which gathers full channel features for predictions, it is evident that substituting the SFC layer with an FC layer yields only marginal improvements at the expense of a much higher storage cost. Furthermore, we compare our method with the model using pooled features for weight prediction in the last row of Table 7. Obviously, without the global fusion, such as the FC or SFC layers, the features extracted from the pointwise convolution network are insufficient to achieve high performance.
@@ 5 Conclusion
This work has proposed the first-of-its-kind purely LUT-based image enhancer, ICELUT, for extremely low-cost and high-speed image retouching. We reveal that input RGB channels are vital for performance and employ fully pointwise convolution kernels that favor subsequent LUT conversion after training. A novel split fully connected layer is devised to effectively suppress the LUT size without compromising performance. With the concerted efforts of these novel designs, our purely LUT-based scheme achieves a remarkable 0.4ms (7ms) on GPU (CPU) with near-state-of-the-art performance, and reduces the power consumption to a negligible level compared to all other CNN solutions.
Acknowledgements
This work was partly supported by the National Natural Science Foundation of China (Grant No. 61991451), the Shenzhen Science and Technology Program (JCYJ20220818101001004), the Theme-Based Research Scheme (TRS) Project T45-701/22-R and General Research Fund (GRF) Project 17203224 of the Research Grants Council (RGC), HKSAR.